# Strong Converse Exponent of Quantum Dichotomies

### v1

@@ Abstract
The quantum dichotomies problem asks at what rate one pair of quantum states can be approximately mapped into another pair of quantum states. In the many copy limit and for vanishing error, the optimal rate is known to be given by the ratio of the respective quantum relative distances. Here, we study the large-deviation behavior of quantum dichotomies and determine the exact strong converse exponent based on the purified distance. Our result is characterized by a simple optimization of quantum Rényi information measures involving all four mutually non-commuting quantum states. Our findings thus constitute an operational example of studying multivariate extensions of quantum Rényi information measure.
@@ I Introduction
Quantum Rényi information measures are basic tools in quantum information theory. Previous measures are usually defined for two non-commuting quantum states, but currently there is an increasing interest in finding quantum Rényi information measure with more than two arguments. The recent works MBV2024geometric ; VTH2024matrix ; FFHT2024matrix have made some progress in this topic (see also BunthVrana2022equivariant ; FLO2023monotonic ; MNW2024on ; NMLW2024optimal ).
One method to find meaningful quantum Rényi information measures is to study the large deviation behaviors of operational quantum information processing tasks. The quantum Rényi information measures usually appear as the large deviation-type exponential rate. In this paper, we study quantum dichotomies transformations. Consider two quantum dichotomies $(\rho_{1},\sigma_{1})$ and $(\rho_{2},\sigma_{2})$, quantum dichotomies transformation is a procedure of using a quantum channel to transform $(\rho_{1},\sigma_{1})$ into as much $(\rho_{2},\sigma_{2})$ as possible, in which we require that the transformation $\sigma_{1}\rightarrow\sigma_{2}$ is exact and the transformation $\rho_{1}\rightarrow\rho_{2}$ is allowed to be approximate FDT2019information ; WangWilde2019resource . It is an important quantum information processing task which have wide variety of applications in quantum resource theory FDT2019information ; WangWilde2019resource ; LPCRTK2024quantum ; KCT2019avoid ; KCT2018Beyond ; KCT2019Moderate ; KumagaiHayashi2013second ; Wilde2022distinguishability . Because there exists four mutually non-commuting quantum states in this problem, it is a proper task for finding multi-variate quantum Rényi information measures. The commonly used distances to measure the closeness between the final state and the target state $\rho_{2}$ are the trace distance and the purified distance. In the asymptotic setting when unlimited $(\rho_{1},\sigma_{1})$ are available, no matter which distance we choose, the optimal rate at which the asymptotically perfect transformation can be achieved equals to the ratio between the quantum relative entropy $D(\rho_{1}\|\sigma_{1})$ and $D(\rho_{2}\|\sigma_{2})$ FDT2019information ; WangWilde2019resource .
Although the first-order asymptotic analysis of quantum dichotomies transformation has been well understood, the finer asymptotic results, including the second-order expansion, moderate-deviation analysis and large-deviation type exponential analysis, are more complicated and they depend on the choice of the distance. The reference LPCRTK2024quantum established the exact second-order asymptotic, moderate-deviation expansion and large-deviation type exponential behavior based on the trace distance under the assumption $\rho_{2}$ commutes with $\sigma_{2}$. The reference KCT2018Beyond and KCT2019Moderate established the second-order asymptotic and the moderate-deviation analysis based on the purified distance, respectively, when $(\rho_{1},\sigma_{1})$ and $(\rho_{2},\sigma_{2})$ are both classical dichotomies. To the best of our knowledge, the current research on the finer asymptotic analysis of quantum dichotomies transformation are only restricted in classical or semi-classical setting. The related results are still unknown in fully quantum setting.
@@
where $D_{\alpha}^{*}(\rho\|\sigma)$ is the sandwiched Rényi divergence. Our work is the first time to establish the exact strong converse exponent for this task in fully quantum setting, filling the gap in previous work. Our strong converse exponent is characterized by a quantum Rényi information measure involving four quantum states.
@@
where $D_{\alpha}(\rho\|\sigma)$ is the Petz Rényi divergence. We conjecture that this formula still holds for general quantum states $\rho_{2}$. The Petz Rényi divergence was used to characterize the reliability functions (the best rate of exponential convergence of the error of a task towards the perfect) for quantum information processing tasks in previous research AKCMBMA2007discriminating ; NussbaumSzkola2009chernoff ; Li2016discriminating ; Nagaoka2006converse ; Hayashi2007error ; ANSV2008asymptotic ; Renes2022achievable ; CHDH2020non ; HayashiTomamichel2016correlation . If the conjecture is correct, this will provide new type of operational meaning for the Petz Rényi divergence.
In the derivation of the achievability part of the strong converse exponent based on the purified distance, we use a technique developed by Mosonyi and Ogawa MosonyiOgawa2017strong , to derive a weaker bound characterized by the log-Euclidean Rényi divergence and then improve it to the final bound, one of the authors also use the similar technique to establish the strong converse exponents for privacy amplification and quantum information decoupling LiYao2024operational . The proof of the optimality part comes from an application of the operator Hölder inequality essentially. For the strong converse exponent based on the trace distance, we mainly make use of an improved Fuchs-van de Graaf to relate it to the strong converse exponent based on the purified distance.
The remainder of this paper is organized as follows. In Section II, we introduce some preliminaries in quantum information theory. In Section III, we present the main problem and the main result. In Section IV, we give the proof of the achievability part of the strong converse exponent for quantum dichotomies transformation. In Section V, the corresponding proof of the optimality part is presented. In Section VI, we conclude this paper with some discussion.
@@ II Preliminaries
II.1 Notation
Let $\mathcal{H}$ be a finite-dimensional Hilbert space, we denote the set of linear operators on $\mathcal{H}$ as $\mathcal{L}(\mathcal{H})$. Let $\mathcal{P}(\mathcal{H})$ and $\mathcal{S}(\mathcal{H})$ represent the set of positive semi-definite operators and the set of quantum states, respectively. The support of an operator $A\in\mathcal{L}(\mathcal{H})$ is denoted as ${\operatorname{supp}}(A)$. For $\rho\in\mathcal{S}(\mathcal{H})$, we use the notation $\mathcal{S}_{\rho}$ for the set of quantum states whose supports are contained in ${\operatorname{supp}}(\rho)$. We denote as $|\mathcal{H}|$ the dimension of $\mathcal{H}$.
@@ II.1 Notation
The pinching inequality Hayashi2002optimal tells that for any $\sigma\in\mathcal{P}(\mathcal{H})$, we have
@@
II.2 Quantum Rényi divergences
The classical Rényi divergence Renyi1961measures is an information quantity which play important roles in many fields. Because of the non-commutativity of quantum states, there are infinite in-equivalent quantum generalizations of classical Rényi divergence. In this section, we introduce the sandwiched Rényi divergence MDSFT2013on ; WWY2014strong , the Petz Rényi divergence Petz1986quasi and the log-Euclidean Rényi divergence HiaiPetz1993the ; MosonyiOgawa2017strong which have been found wide applications so far AKCMBMA2007discriminating ; NussbaumSzkola2009chernoff ; Li2016discriminating ; Nagaoka2006converse ; Hayashi2007error ; ANSV2008asymptotic ; Renes2022achievable ; CHDH2020non ; HayashiTomamichel2016correlation ; MosonyiOgawa2015quantum ; GuptaWilde2015multiplicativity ; CMW2016strong ; LiYao2024strong ; LYH2023tight ; LiYao2024reliability ; Hayashi2015precise ; LiYao2021reliable .
Definition 1
Let $\alpha\in(0,+\infty)\setminus\{1\}$, $\rho\in\mathcal{S}(\mathcal{H})$ and $\sigma\in\mathcal{P}(\mathcal{H})$. When $\alpha>1$ and ${\operatorname{supp}}(\rho)\subseteq{\operatorname{supp}}(\sigma)$ or $\alpha\in(0,1)$ and ${\operatorname{supp}}(\rho)\not\perp{\operatorname{supp}}(\sigma)$, the sandwiched Rényi divergence and the Petz Rényi divergence of order $\alpha$ are respectively defined as
@@ II.2 Quantum Rényi divergences
otherwise, we set $D_{\alpha}^{*}(\rho\|\sigma)=D(\rho\|\sigma)=+\infty$. When $\alpha>1$ and ${\operatorname{supp}}(\rho)\subseteq{\operatorname{supp}}(\sigma)$ or $\alpha\in(0,1)$ and ${\operatorname{supp}}(\rho)\cap{\operatorname{supp}}(\sigma)\neq\{0\}$, the log-Euclidean Rényi divergence of order $\alpha$ is defined as
@@
When $\alpha\rightarrow 1$, all above quantum Rényi divergences converge to the Umegaki relative entropy Umegaki1954conditional
$D(\rho\|\sigma):=\begin{cases}\operatorname{Tr}\rho(\log\rho-\log\sigma)&\text{ if }{\operatorname{supp}}(\rho)\subseteq{\operatorname{supp}}(\sigma),\\ +\infty&\text{ otherwise.}\end{cases}$ (7)
@@
    Monotonicity in Rényi parameter MDSFT2013on ; Beigi2013sandwiched ; MosonyiOgawa2017strong : if $0\leq\alpha\leq\beta$, then $D_{\alpha}^{(t)}(\rho\|\sigma)\leq D^{(t)}_{\beta}(\rho\|\sigma)$, for $(t)=*$, $(t)=\{\}$ and $(t)=\flat$;
@@
    Monotonicity in $\sigma$ MDSFT2013on ; MosonyiOgawa2017strong : if $\sigma^{\prime}\geq\sigma$, then $D_{\alpha}^{(t)}(\rho\|\sigma^{\prime})\leq D_{\alpha}^{(t)}(\rho\|\sigma)$, for $(t)=*$, $\alpha\in[\frac{1}{2},+\infty)$, for $(t)=\{\}$, $\alpha\in[0,2]$ and for $(t)=\flat$, $\alpha\in[0,+\infty)$;
@@
    Variational representation MosonyiOgawa2017strong : the log-Euclidean Rényi divergence has the following variational representation
@@
    Data processing inequality FrankLieb2013monotonicity ; Beigi2013sandwiched ; MDSFT2013on ; WWY2014strong ; MosonyiOgawa2017strong : letting $\mathcal{N}$ be a CPTP map from $\mathcal{L}(\mathcal{H})$ to $\mathcal{L}(\mathcal{H}^{\prime})$, we have
@@
    Approximation by pinching HayashiTomamichel2016correlation ; MosonyiOgawa2015quantum : for any $\alpha\geq 0$, we have
@@ III Problem formulation and main result
where $\Delta$ can be chosen as the trace distance or the purified distance.
In the asymptotic setting in which multiple copies of $(\rho_{1},\sigma_{1})$ are available, the first order asymptotic behavior of $M^{\Delta,\epsilon}_{\rho_{1},\sigma_{1}\rightarrow\rho_{2},\sigma_{2}}(n)$ has been derived in FDT2019information ; WangWilde2019resource , i.e.,
@@
where $\Delta$ can be either trace distance or purified distance.
@@
In this paper, we derive the exact expression for $E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$ and if $\rho_{2}$ is a pure state, we also determine the exact expression for $E^{d}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$. Our main results are as follows.
Theorem 5
For any $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ satisfy ${\operatorname{supp}}(\rho_{1})\subseteq{\operatorname{supp}}(\sigma_{1})$ or ${\operatorname{supp}}(\rho_{2})\subseteq{\operatorname{supp}}(\sigma_{2})$ and any $r>0$, we have
@@
According to the definition of $E^{d}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$ and $E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$, we have
@@ IV Proof of the achievability part
$F(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r):=\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{\flat}(\rho_{2}\|\sigma_{2})-D_{\beta}^{\flat}(\rho_{1}\|\sigma_{1})\big{\}},$ (19)
@@
Proposition 8
For $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$ that satisfy ${\operatorname{supp}}(\rho_{1})\subset{\operatorname{supp}}(\sigma_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ and $r\geq 0$, we have
$F(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)=\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\{rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+|rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})|^{+}\}.$
Proof. By Proposition 2(iii), we have
$\begin{split}&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{\flat}(\rho_{2}\|\sigma_{2})-D_{\beta}^{\flat}(\rho_{1}\|\sigma_{1})\big{\}}\\ =&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}r\inf_{\tau_{2}\in\mathcal{S}_{\rho_{2}}}\big{(}D(\tau_{2}\|\sigma_{2})-\frac{\alpha}{\alpha-1}D(\tau_{2}\|\rho_{2})\big{)}-\sup_{\tau_{1}\in\mathcal{S}_{\rho_{1}}}\big{(}D(\tau_{1}\|\sigma_{1})-\frac{\beta}{\beta-1}D(\tau_{1}\|\rho_{1})\big{)}\big{\}}\\ =&\sup_{\frac{1}{2}<\alpha<1}\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\frac{1-\alpha}{\alpha}\big{\{}rD(\tau_{2}\|\sigma_{2})-\frac{\alpha}{\alpha-1}rD(\tau_{2}\|\rho_{2})-D(\tau_{1}\|\sigma_{1})+\frac{\beta}{\beta-1}D(\tau_{1}\|\rho_{1})\big{\}}\\ =&\sup_{\frac{1}{2}<\alpha<1}\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\big{\{}rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+\frac{1-\alpha}{\alpha}\big{(}rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})\big{)}\big{\}}\\ =&\sup_{0<\delta<1}\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\big{\{}rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+\delta\big{(}rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})\big{)}\big{\}}\\ =&\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\sup_{0<\delta<1}\big{\{}rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+\delta\big{(}rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})\big{)}\big{\}}\\ =&\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\{rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+|rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})|^{+}\},\end{split}$ (20)
where the sixth line follows from Sion’s minimax theorem. The convexity of the function $(\tau_{1},\tau_{2})\mapsto rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+\delta\big{(}rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})$ that the Sion’s minimax theorem requires is not obvious. To see this, we write
@@
[…]
@@
Lemma 9
For $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$ satisfy ${\operatorname{supp}}(\rho_{1})\subset{\operatorname{supp}}(\sigma_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ and $r\geq 0$, we have
@@
$\begin{split}&F_{1}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r):=\inf_{(\tau_{1},\tau_{2})\in\mathcal{F}_{1}}\{rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})\}\\ &F_{2}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r):=\inf_{(\tau_{1},\tau_{2})\in\mathcal{F}_{2}}\{rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})\},\end{split}$ (22)

### v2

@@ Abstract
The quantum dichotomies problem asks at what rate one pair of quantum states can be approximately mapped into another pair of quantum states. In the many copy limit and for vanishing error, the optimal rate is known to be given by the ratio of the respective quantum relative distances. Here, we study the large-deviation behavior of quantum dichotomies and determine the exact strong converse exponent based on the purified distance. This is the first time to establish the exact high-error large-deviation analysis for this task in fully quantum setting.
Index Terms:
strong converse exponent, quantum dichotomies, sandwiched Rényi divergence, purified distance, trace distance
@@ I Introduction
Consider two quantum dichotomies $(\rho_{1},\sigma_{1})$ and $(\rho_{2},\sigma_{2})$, quantum dichotomies transformation is a procedure of using a quantum channel to transform $(\rho_{1},\sigma_{1})$ into as much $(\rho_{2},\sigma_{2})$ as possible, in which we require that the transformation $\sigma_{1}\rightarrow\sigma_{2}$ is exact and the transformation $\rho_{1}\rightarrow\rho_{2}$ is allowed to be approximate [1, 2]. It is an important quantum information processing task which have wide variety of applications in quantum resource theory [1, 2, 3, 4, 5, 6, 7, 8]. The commonly used distances to measure the closeness between the final state and the target state $\rho_{2}$ are the trace distance and the purified distance. In the asymptotic setting when unlimited $(\rho_{1},\sigma_{1})$ are available, no matter which distance we choose, the optimal rate at which the asymptotically perfect transformation can be achieved equals to the ratio between the quantum relative entropy $D(\rho_{1}\|\sigma_{1})$ and $D(\rho_{2}\|\sigma_{2})$ [1, 2].
Although the first-order asymptotic analysis of quantum dichotomies transformation has been well understood, the finer asymptotic results, including the second-order expansion, moderate-deviation analysis and large-deviation type exponential analysis, are more complicated and they depend on the choice of the distance. The reference [3] established the exact second-order asymptotic, moderate-deviation expansion and large-deviation type exponential behavior based on the trace distance under the assumption $\rho_{2}$ commutes with $\sigma_{2}$. The reference [5] and [6] established the second-order asymptotic and the moderate-deviation analysis based on the purified distance, respectively, when $(\rho_{1},\sigma_{1})$ and $(\rho_{2},\sigma_{2})$ are both classical dichotomies. To the best of our knowledge, the current research on the finer asymptotic analysis of quantum dichotomies transformation are only restricted in classical or semi-classical setting. The related results are still unknown in fully quantum setting.
@@
where $D_{\alpha}^{*}(\rho\|\sigma)$ is the sandwiched Rényi divergence. Our work is the first time to establish the exact strong converse exponent for this task in fully quantum setting, filling the gap in previous work.
@@
where $D_{\alpha}(\rho\|\sigma)$ is the Petz Rényi divergence. We conjecture that this formula still holds for general quantum states $\rho_{2}$. The Petz Rényi divergence was used to characterize the reliability functions (the best rate of exponential convergence of the error of a task towards the perfect) for quantum information processing tasks in previous research [9, 10, 11, 12, 13, 14, 15, 16, 17]. If the conjecture is correct, this will provide new type of operational meaning for the Petz Rényi divergence.
In the derivation of the achievability part of the strong converse exponent based on the purified distance, we use a technique developed by Mosonyi and Ogawa [18], to derive a weaker bound characterized by the log-Euclidean Rényi divergence and then improve it to the final bound, one of the authors also use the similar technique to establish the strong converse exponents for privacy amplification and quantum information decoupling [19]. The proof of the optimality part comes from an application of the operator Hölder inequality and this part can also be deduced from the work of Wang and Wilde [2]. For the strong converse exponent based on the trace distance, we mainly make use of an improved Fuchs-van de Graaf to relate it to the strong converse exponent based on the purified distance.
The remainder of this paper is organized as follows. In Section II, we introduce some preliminaries in quantum information theory. In Section III, we present the main problem and the main result. In Section IV, we give the proof of the achievability part of the strong converse exponent for quantum dichotomies transformation. In Section V, we conclude this paper with some discussion.
@@ II Preliminaries
II-A Notation
Let $\mathcal{H}$ be a finite-dimensional Hilbert space, we denote the set of linear operators on $\mathcal{H}$ as $\mathcal{L}(\mathcal{H})$. Let $\mathcal{P}(\mathcal{H})$ and $\mathcal{S}(\mathcal{H})$ represent the set of positive semi-definite operators and the set of quantum states, respectively. The support of an operator $A\in\mathcal{L}(\mathcal{H})$ is denoted as $\mathop{\rm supp}(A)$. For $\rho\in\mathcal{S}(\mathcal{H})$, we use the notation $\mathcal{S}_{\rho}$ for the set of quantum states whose supports are contained in $\mathop{\rm supp}(\rho)$. We denote as $|\mathcal{H}|$ the dimension of $\mathcal{H}$.
@@ II-A Notation
The pinching inequality [20] tells that for any $\sigma\in\mathcal{P}(\mathcal{H})$, we have
@@
II-B Quantum Rényi divergences
The classical Rényi divergence [21] is an information quantity which play important roles in many fields. Because of the non-commutativity of quantum states, there are infinite in-equivalent quantum generalizations of classical Rényi divergence. In this section, we introduce the sandwiched Rényi divergence [22, 23], the Petz Rényi divergence [24] and the log-Euclidean Rényi divergence [18, 25] which have been found wide applications so far [9, 10, 11, 12, 13, 14, 15, 16, 17, 26, 27, 28, 29, 30, 31, 32, 33].
Definition 1.
Let $\alpha\in(0,+\infty)\setminus\{1\}$, $\rho\in\mathcal{S}(\mathcal{H})$ and $\sigma\in\mathcal{P}(\mathcal{H})$. When $\alpha>1$ and $\mathop{\rm supp}(\rho)\subseteq\mathop{\rm supp}(\sigma)$ or $\alpha\in(0,1)$ and $\mathop{\rm supp}(\rho)\not\perp\mathop{\rm supp}(\sigma)$, the sandwiched Rényi divergence and the Petz Rényi divergence of order $\alpha$ are respectively defined as
@@ II-B Quantum Rényi divergences
otherwise, we set $D_{\alpha}^{*}(\rho\|\sigma)=D(\rho\|\sigma)=+\infty$. When $\alpha>1$ and $\mathop{\rm supp}(\rho)\subseteq\mathop{\rm supp}(\sigma)$ or $\alpha\in(0,1)$ and $\mathop{\rm supp}(\rho)\cap\mathop{\rm supp}(\sigma)\neq\{0\}$, the log-Euclidean Rényi divergence of order $\alpha$ is defined as
@@
When $\alpha\rightarrow 1$, all above quantum Rényi divergences converge to the Umegaki relative entropy [34]
$D(\rho\|\sigma):=\begin{cases}\operatorname{Tr}\rho(\log\rho-\log\sigma)&\text{ if }\mathop{\rm supp}(\rho)\subseteq\mathop{\rm supp}(\sigma),\\ +\infty&\text{ otherwise.}\end{cases}$ (7)
@@
    Monotonicity in Rényi parameter [18, 22, 35]: if $0\leq\alpha\leq\beta$, then $D_{\alpha}^{(t)}(\rho\|\sigma)\leq D^{(t)}_{\beta}(\rho\|\sigma)$, for $(t)=*$, $(t)=\{\}$ and $(t)=\flat$;
@@
    Monotonicity in $\sigma$ [18, 22]: if $\sigma^{\prime}\geq\sigma$, then $D_{\alpha}^{(t)}(\rho\|\sigma^{\prime})\leq D_{\alpha}^{(t)}(\rho\|\sigma)$, for $(t)=*$, $\alpha\in[\frac{1}{2},+\infty)$, for $(t)=\{\}$, $\alpha\in[0,2]$ and for $(t)=\flat$, $\alpha\in[0,+\infty)$;
@@
    Variational representation [18]: the log-Euclidean Rényi divergence has the following variational representation
@@
    Data processing inequality [18, 22, 23, 35, 36]: letting $\mathcal{N}$ be a CPTP map from $\mathcal{L}(\mathcal{H})$ to $\mathcal{L}(\mathcal{H}^{\prime})$, we have
@@
    Approximation by pinching [17, 26]: for any $\alpha\geq 0$, we have
@@ III Problem formulation and main result
where $\Delta$ can be chosen as the trace distance $d$ or the purified distance $P$.
In the asymptotic setting in which multiple copies of $(\rho_{1},\sigma_{1})$ are available, the first order asymptotic behavior of $M^{\Delta,\epsilon}_{\rho_{1},\sigma_{1}\rightarrow\rho_{2},\sigma_{2}}(n)$ has been derived in [1, 2] , i.e.,
@@
where $\Delta$ can be either $d$ or $P$.
@@
Wang and Wilde have established a lower bound for $E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$ in [2, Proposition 1], following their discussion in Eq. (J33)–(J35) and (L11)–(J12). In particular, the following lower bound can be derived
$E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)\geq\sup_{\frac{1}{2}\leq\alpha\leq 1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\frac{\alpha}{2\alpha-1}}^{*}(\rho_{1}\|\sigma_{1})\big{\}}.$ (14)
In this paper, we prove that the lower bound in Eq. (14) is also achievable, thus establishing the exact expression for $E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$ and if $\rho_{2}$ is a pure state, we also determine the exact expression for $E^{d}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$. Our main results are as follows.
Theorem 4.
For any $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ that satisfy $\mathop{\rm supp}(\rho_{1})\subseteq\mathop{\rm supp}(\sigma_{1})$ or $\mathop{\rm supp}(\rho_{2})\subseteq\mathop{\rm supp}(\sigma_{2})$ and any $r>0$, we have
@@
If Eq. (15) holds, we can obtain from Eq. (18)
@@ IV Proof of the achievability part
$F(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r):=\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\left\{rD_{\alpha}^{\flat}(\rho_{2}\|\sigma_{2})-D_{\beta}^{\flat}(\rho_{1}\|\sigma_{1})\right\},$ (20)
@@
Proposition 6.
For $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$ that satisfy $\mathop{\rm supp}(\rho_{1})\subset\mathop{\rm supp}(\sigma_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ and $r\geq 0$, we have
$F(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)=\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\left\{rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+|rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})|^{+}\right\}.$
Proof: By Proposition 2(iii), we have
$\begin{split}&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\{rD_{\alpha}^{\flat}(\rho_{2}\|\sigma_{2})-D_{\beta}^{\flat}(\rho_{1}\|\sigma_{1})\}\\ =&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}r\inf_{\tau_{2}\in\mathcal{S}_{\rho_{2}}}\big{(}D(\tau_{2}\|\sigma_{2})-\frac{\alpha}{\alpha-1}D(\tau_{2}\|\rho_{2})\big{)}-\sup_{\tau_{1}\in\mathcal{S}_{\rho_{1}}}\big{(}D(\tau_{1}\|\sigma_{1})-\frac{\beta}{\beta-1}D(\tau_{1}\|\rho_{1})\big{)}\big{\}}\\ =&\sup_{\frac{1}{2}<\alpha<1}\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\frac{1-\alpha}{\alpha}\{rD(\tau_{2}\|\sigma_{2})-\frac{\alpha}{\alpha-1}rD(\tau_{2}\|\rho_{2})-D(\tau_{1}\|\sigma_{1})+\frac{\beta}{\beta-1}D(\tau_{1}\|\rho_{1})\}\\ =&\sup_{\frac{1}{2}<\alpha<1}\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\big{\{}rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+\frac{1-\alpha}{\alpha}\big{(}rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})\big{)}\big{\}}\\ =&\sup_{0<\delta<1}\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\big{\{}rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+\delta\big{(}rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})\big{)}\big{\}}\\ =&\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\sup_{0<\delta<1}\big{\{}rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+\delta\big{(}rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})\big{)}\big{\}}\\ =&\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\big{\{}rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+|rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})|^{+}\big{\}},\end{split}$ (21)
where the sixth line follows from Sion’s minimax theorem. The convexity of the function $(\tau_{1},\tau_{2})\mapsto rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+\delta\big{(}rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})\big{)}$ that the Sion’s minimax theorem requires is not obvious. To see this, we write
@@
Because $rD(\tau_{2}\|\rho_{2})+\delta rD(\tau_{2}\|\sigma_{2})$, $-(1-\delta)H(\tau_{1})$ and $\operatorname{Tr}\tau_{1}(\delta\log\sigma_{1}-\log\rho_{1})$ are all convex function of $(\tau_{1},\tau_{2})$, the result follows.
@@
Lemma 7.
For $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$ that satisfy $\mathop{\rm supp}(\rho_{1})\subset\mathop{\rm supp}(\sigma_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ and $r\geq 0$, we have
@@
$\begin{split}&F_{1}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r):=\inf_{(\tau_{1},\tau_{2})\in\mathcal{F}_{1}}\left\{rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})\right\}\\ &F_{2}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r):=\inf_{(\tau_{1},\tau_{2})\in\mathcal{F}_{2}}\left\{rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})\right\},\end{split}$ (23)